An object of unknown mass m is hung from a vertical spring of unknown spring constant k, and the object is observed to be at rest when the spring has stretch by 14 cm. The object is then given a slight push upward and executes SHM. Determine the period T of this oscillation.

1
Expert's answer
2022-04-11T12:06:01-0400

The period T of the SHM

"T=2\pi\sqrt{\frac{m}{k}}"

Since,

"mg=k\Delta l"

so

"\frac{m}{k}=\frac{\Delta l}{g}"

Finally

"T=2\pi\sqrt{\frac{\Delta l}{g}}=6.28*\sqrt{\frac{0.14}{9.8}}=0.75\:\rm s"